OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3)

OpenStudy (anonymous):

x^2 + 2?

OpenStudy (anonymous):

Oh yea x^2

OpenStudy (anonymous):

3^2 + 2 = 11

OpenStudy (anonymous):

keep forgetting to put that for x squared

OpenStudy (anonymous):

g(x) = x2 + 2 by putting g(3) g(3)=(3^2)+2 g(3)=9+2 g(3)=11